A block sliding to the right along a horizontal frictional surface has an initial velocity of 12.4 m/s and stops in 110 m due to friction with the surface.

Find the acceleration of the block.

Find the coefficient of kinetic friction between the block and the surface.
Answer #1

u =12.4 m/s , v =0, s =110 m

From kinematic equation

v^2 -u^ =2as

0 -u^2 =2as

a = -u^2/2s = -[12.4*12.4]/(2*110)

a = -0.699 m/s2

Fk = ma

uk *N= ma

uk*mg =ma

uk =a/g = 0.699/9.8

uk = 0.0713